OpenStudy (anonymous):

write an equation of the line with slope 3 and goes through the points (1,4)

OpenStudy (triciaal):

points (x, y) and (1, 4) slope of line = change in y divided by change in x given = 3 m = (y2- y1)/(x2 - x1) m is the slope and you have the values for 2 points

OpenStudy (anonymous):

but whats the answer?:b @triciaal

jimthompson5910 (jim_thompson5910):

there's no need for m = (y2- y1)/(x2 - x1) since they give you the slope already

OpenStudy (anonymous):

i need help working through this problem so i know im doing it right @jim_thompson5910

jimthompson5910 (jim_thompson5910):

use y = mx+b

jimthompson5910 (jim_thompson5910):

m = 3 is the given slope (x,y) = (1,4) is the given point

jimthompson5910 (jim_thompson5910):

use those pieces of info to find b

OpenStudy (triciaal):